When an ac source with a peak potential difference of 42V is connected to a series RLC circuit, the peak current is 4 A. The capacitor is C = 20 μF, the frequency is 60 Hz, and the current leads the potential difference by 30 ∘ degrees. Find the resitance and the inductance.

R = Ω

L = H
Answer #1

R L womt Vm =42v Impedance 2= R + j(XL-Xc) =Rt j (CWL- uc), R²+ (WL- wc)2. 121- Im = peak current - 4 Amp Im = Vm 42 121 ✓ R²into ean Now put w Lotuc = 2 R² + ( R ) 441 4 R + R? 3 441 4 → 4 R² - 4211 = R² = 44183 46x4 44103 4X4 = 2113 =) R = 4. W = 2